OpenStudy (anonymous):

HELPPP GEOMETRY UNIT2 LESSON% MEDAL DOE BEST ANSWER!!!

OpenStudy (anonymous):

@Joe711

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

D:

OpenStudy (anonymous):

Is it multiple choice or?

OpenStudy (anonymous):

no D:

OpenStudy (anonymous):

i reallt need help

OpenStudy (anonymous):

SOH CAH TOA \[\sin \theta =\frac{ Opposite }{Hypoteneuse } \] \[\cos \theta =\frac{ adjacent}{Hypoteneuse } \] \[\tan\theta =\frac{ adjacent}{Hypoteneuse } \] We know that

OpenStudy (anonymous):

oops... for tan...it is\[\tan \theta =\frac{ opposite}{adjacent } \] my bad

OpenStudy (anonymous):

you have to use sin here because 500m is the Opposite of theta=10deg... x is the Hypotenuese